OpenStudy (anonymous):

5x - 2y = -1 x + 4y = 35 solve using the substitution method & then using the addition method

OpenStudy (anonymous):

x=3, y=8

OpenStudy (anonymous):

substitution method: 5x-2y=-1 x=35-4y (subtracted 4y) (substitute second equation into first) 5(35-4y)-2y=-1 175-20y-2y=-1 (multiplied the 5*(35-4y) -20y-2y=-176 (subtracted 175) -22y=-176 (-2y+-20y) y=8 (divided by 22) now just put into second equation x=35-4(8) x=35-32 x=3

OpenStudy (anonymous):

addition method: 5x-2y=-1 x+ 4y=35 10x-4y=-2 (multiplied first equation by 2) x+4y=35 11x+0y=33 (added 2 equations) 11x=33 x=3 x+4y=35 3+4y=35 (substitute x=3) 4y=32 (subtract 3) y=8

OpenStudy (anonymous):

ok, i'm looking it all over to seer if i understand it! thank u!